(> Le* I S A is alinear transfor mation on R* (in the standard bosis vasiS s is another basis for R Prololem: Rewrite the wottix A in the basis ^
Answer #1

Given, A = 045 211 130 is a linear transformation on R3 in the standard basis ,i.e., \left \{ \begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix},\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix},\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix} \right \} .

Now, T\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix} = 1\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix}+3\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}+0\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix} = \begin{bmatrix} 1\\ 3\\ 0 \end{bmatrix}

And, T\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix} = 2\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix}+1\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}+1\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix} = \begin{bmatrix} 2\\ 1\\ 1 \end{bmatrix}

And, T\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix} = 0\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix}+4\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}+5\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix} = \begin{bmatrix} 0\\ 4\\ 5 \end{bmatrix}

Let \begin{bmatrix} x\\ y\\ z \end{bmatrix} \in R3.

Then, \begin{bmatrix} x\\ y\\ z \end{bmatrix} = x\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix}+y\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}+z\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix}

i.e., T\begin{bmatrix} x\\ y\\ z \end{bmatrix} = x*T\begin{bmatrix} 1\\ 0\\ 0 \end{bmatrix}+y*T\begin{bmatrix} 0\\ 1\\ 0 \end{bmatrix}+z*T\begin{bmatrix} 0\\ 0\\ 1 \end{bmatrix}

i.e., T\begin{bmatrix} x\\ y\\ z \end{bmatrix} = x\begin{bmatrix} 1\\ 3\\ 0 \end{bmatrix}+y\begin{bmatrix} 2\\ 1\\ 1 \end{bmatrix}+z\begin{bmatrix} 0\\ 4\\ 5 \end{bmatrix}

i.e., T\begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} x+2y\\ 3x+y+4z\\ y+5z \end{bmatrix}

Now, T\begin{bmatrix} 1\\ 1\\ 1 \end{bmatrix} = \begin{bmatrix} 1+2\\ 3+1+4\\ 1+5 \end{bmatrix} = \begin{bmatrix} 3\\ 8\\ 6 \end{bmatrix}

And, T\begin{bmatrix} 1\\ 2\\ 2 \end{bmatrix} = \begin{bmatrix} 1+4\\ 3+2+8\\ 2+10 \end{bmatrix} = \begin{bmatrix} 5\\ 13\\ 12 \end{bmatrix}

And, T\begin{bmatrix} 1\\ 2\\ 3 \end{bmatrix} = \begin{bmatrix} 1+4\\ 3+2+12\\ 2+15 \end{bmatrix} = \begin{bmatrix} 5\\ 17\\ 17 \end{bmatrix}

Here, \begin{bmatrix} 3\\ 8\\ 6 \end{bmatrix} = (-2)\begin{bmatrix} 1\\ 1\\ 1 \end{bmatrix}+7\begin{bmatrix} 1\\ 2\\ 2 \end{bmatrix}+(-2)\begin{bmatrix} 1\\ 2\\ 3 \end{bmatrix}

And, \begin{bmatrix} 5\\ 13\\ 12 \end{bmatrix} = (-3)\begin{bmatrix} 1\\ 1\\ 1 \end{bmatrix}+9\begin{bmatrix} 1\\ 2\\ 2 \end{bmatrix}+(-1)\begin{bmatrix} 1\\ 2\\ 3 \end{bmatrix}

And, \begin{bmatrix} 5\\ 17\\ 17 \end{bmatrix} = (-7)\begin{bmatrix} 1\\ 1\\ 1 \end{bmatrix}+12\begin{bmatrix} 1\\ 2\\ 2 \end{bmatrix}+0\begin{bmatrix} 1\\ 2\\ 3 \end{bmatrix}

Therefore, the matrix A in the basis B is = \begin{bmatrix} -2 & -3 & -7\\ 7 & 9 & 12\\ -2 & -1 & 0 \end{bmatrix} .